OpenStudy (anonymous):

What is the value of the expression x-6 over 4 when x = –26?

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

substitute the value of x into\[\frac{ x-6 }{ 4 }\]

OpenStudy (anonymous):

it will be\[\frac{ -26-6 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ -32 }{ 4 }\]

OpenStudy (anonymous):

then -32 divided by 4

OpenStudy (anonymous):

and what would you get?/

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yayyy